# Binary and Non-Binary Self-Dual Sequences and Maximum Period Single-Track Gray Codes

### v1

@@ Abstract
Binary self-dual sequences have been considered and analyzed throughout the years, and they were used for various applications. Motivated by a construction for single-track Gray codes, we examine the structure and recursive constructions for binary and non-binary self-dual sequences. The feedback shift registers that generate such sequences are discussed. The connections between these sequences and maximum period single-track codes are discussed. Maximum period non-binary single-track Gray codes of length $p^{t}$ and period $p^{p^{t}}$ are constructed. These are the first infinite families of maximum period codes presented in the literature.
@@ I Introduction
Motivated by these applications and mainly by single-track Gray codes, we are trying to find properties of these sequences, to construct them recursively and efficiently, and to order them for the construction of single-track Gray codes. The Gray code, which we consider, will be only for maximum period codes. This will restrict some of the parameters of the self-dual sequences. We consider the generalization of such sequences over a non-binary alphabet, where a non-binary self-dual sequence is a sequence that is invariant under the addition of a nonzero constant to all the elements of the sequence.
@@ II Preliminaries
The consecutive values obtained from $x_{n+1}$ (until the initial state is repeated), starting with any initial state is the sequence generated by the register. The CCRn generates all the SDSs whose length is a divisor of $2n$ and no other sequences. The period of a sequence $S=[s_{1}s_{2}~\cdots~s_{k}]$ is the length of the shortest word $X$ for which $S=[X,X,\ldots,X]$. If the FSRn is linear, then it can also be represented by a polynomial of degree $n$ over the binary field.
The number of sequences of the CCRn was presented in [11].
@@
Let SD(n)(n) denote the number of SDSs of period $n$. The following connection between the number of sequences generated by the CCRn and the number of SDSs of various periods was proved in [6].
@@
where $Z$ can be any sequence of length $2^{n}$. There are $2^{2^{n}-1}$ distinct cyclic sequences generated in this way (if $Z$ is replaced by $Z+X+Y$ the same sequence is obtained). If $Y=\bar{X}$ then
@@
It was proved in [16] that there is no maximum period STGC of length $2^{t}$ and period $2^{2^{t}}$, $t>1$. An STGC of length $2^{t}$ and period $2^{2^{t}}-2^{t+1}$ was constructed in [9]. Non-binary STGCs of length $p^{m}$ and period $p^{p^{m}}$ are discussed in Section V. These are the only possible maximum period STGCs that contain all the words of the given length. For each odd prime $p$, we might have a maximum period STGC of length $p$ and period $2^{p}-2$. Such codes are known to exist for all primes up to $p=19$ [4], and they are constructed using the construction of Theorem 3, but they can also be obtained by the construction of Theorem 4. Maximum period STGC using these theorems can also be obtained for $n=9$ and $n=25$. For other values of $n$, the construction of Theorem 3 cannot yield a maximum period STGCs. The situation for the construction of Theorem 4 is quite different. If $n=2p$, then there is exactly one sequence of period 4 and $\frac{2^{2p}-4}{4p}$ sequences of period $4p$ generated by the CCR2p. It might be possible, in this case, to use the construction of Theorem 4 to generate a maximum period STGC. If $n=4p$, then there are exactly two sequences of period 8 and $\frac{2^{4p}-16}{8p}$ sequences of period $8p$ generated by the CCR4p. It might be possible, in this case, to use the construction of Theorem 4 to generate a maximum period STGC. If $n=8p$, then there are exactly sixteen sequences of period 16 and $\frac{2^{8p}-256}{16p}$ sequences of period $16p$ generated by the CCR8p and it might be possible to use the construction of Theorem 4 to generate a maximum period STGC when $p>16$. The same process can be analyzed for $n=2^{i}p$, $i\geq 1$.
@@ III Binary Self-Dual Sequences
The operation ${\bf D}^{2^{n}}=({\bf E}+1)^{2^{n}}$ and its inverse are associated with the sequences generated by the polynomial $(x+1)^{2^{n}}$. They are generalized for binary sequences whose length is not a power of 2 in the following way. Let ${\bf\Delta}_{m}$ be the following operator defined on binary sequences whose length is $rm$. Let $S=[X_{1},X_{2},\ldots,X_{r}]$ be a binary sequence, where $X_{i}$, $1\leq i\leq r$, is a word of length of length $m$. The operator ${\bf\Delta}_{m}$ is defined on $S$ as (note that plus and minus are the same in the binary field)
@@
If $\cal S$} is a set of sequences, then we define ${\bf\Delta}_{m}{\cal S}$ to be the set $\{{\bf\Delta}_{m}T~:~T\in{\cal S}\}$. and ${\bf\Delta}_{m}^{-1}{\cal S}$ is the set $\{{\bf\Delta}_{m}^{-1}T~:~T\in{\cal S}\}$. It is readily verified that for any sequence $S$ of length $rm$ we have that ${\bf\Delta}_{m}({\bf\Delta}_{m}^{-1}S)=S$.
@@
An SDS of length $2n$ generated by the CCRn can be written in the form $X,\bar{X}]$, where the length of $X$ is $n$. Hence, ${\bf\Delta}_{n}^{-1}[X,\bar{X}]=[Y,X+Y,\bar{Y},X+\bar{Y}]$, where $Y$ is any sequence of length $n$, has length $4n$ and it is an SDS generated by the CCR2n. ∎
@@
There are two SDSs generated by the CCR3, $[000111]$ and $[01]$. By applying ${\bf\Delta}_{3}^{-1}$ on these two sequences, the six sequences of the CCR6 are obtained. Four sequences of period 12
@@
are obtained by applying ${\bf\Delta}_{3}^{-1}$ on $[000111]$. One sequence of period 12, [000010111101][000~~010~~111~~101], and one sequence of period 4, $[0011]$, are obtained by applying ${\bf\Delta}_{3}^{-1}$ on $[01]$.
We can continue and apply ${\bf\Delta}_{6}^{-1}$ on these 6 SDSs of the CCR6. By applying ${\bf\Delta}_{6}^{-1}$ on the five sequences of period 12, we obtain 160 SDSs of period 24, which are generated by the CCR12. By applying ${\bf\Delta}_{6}^{-1}$ on the sequence $[0011]$ we obtain the two SDSs $[00001111]$ and $[00101101]$ of period 8 and ten SDSs of period 24, all generated by the CCR12.
@@
The recurrence of Theorem 5 and the enumeration of Theorem 6 and Example 1 can be further generalized and analyzed. From the $\frac{2^{2^{i+1}p}-2^{2^{i+1}}}{2^{i+2}p}$ SDSs of period $2^{i+2}p$ only $\frac{2^{2^{i}(p+1)}-2^{2^{i+1}}}{2^{i+2}p}$ are generated from SDSs of period $2^{i+1}$, where each such sequence of period $2^{i+2}p$ can be generated using any one of $p$ different words of length $2^{i}p-1$. It can be analyzed which words of length $2^{i}p-1$ yield SDSs of period $2^{i+2}$.
@@ IV Non-binary Self-Dual Sequences
Non-binary self-dual sequences were hardly considered. We generalize the definition of the CCRn for an FSRn over $\mathbb{Z}_{m}$. The $m$-CCRn is an FSRn over $\mathbb{Z}_{m}$ with feedback function
@@
where $x_{i}\in\mathbb{Z}_{m}$, $1\leq i\leq n$, and the computation is performed modulo $m$. The function can also be taken as $f(x_{1},x_{2},\ldots,x_{n})=x_{1}+r$, where $r$ is relatively prime to $m$, but it will not affect the properties of the generated sequences. The enumerations of the number of sequences of the CCRn can be generalized. For example, we have the following generalizations of Theorem 1 and Lemma 1.
@@
The period of a binary SDS generated by the CCRn is any integer $d$, where $d$ divides $2n$ but does not divide $n$. This is generalized for the $m$-CCRn as follows.
@@
$\sum_{\begin{subarray}{c}d|mn\\ d\nmid n\\ d|n-\frac{d}{m}\end{subarray}}\textup{SD}(d).$
@@
The sequence $V,V+{\bf 1},V+{\bf 2}]$ (a bold integer is a sequence with only this integer), where $V=(Z,Z+Y,Z+2Y+X)$, is an SDS of period $3^{n+1}$ over $\mathbb{Z}_{3}$. Each different choice of $Z$ and $Y$ yields a distinct SDS. Each SDS sequence of period $3^{n+1}$ over $\mathbb{Z}_{3}$ is constructed in this way.
@@
The proof has similarity to that of Theorem 5 with one major exception. By the definition of ${\bf D}^{-1}$ we have that D^-3^n-1 [X , X+1, X+2] = [Y,Y+X, Y+2X+ 1] and D^-3^n-1 [Y,Y+X, Y+2X+ 1] = [V , V+1, V+2], where $V=(Z,Z+Y,Z+2Y+X)$. Note, that since either $V$ or $V+{\bf 1}$ or $V+{\bf 2}$ starts with a zero, it was sufficient to consider for $Z$ only those words of length $3^{n-1}$ that start with a zero.
@@
This implies that each SDS sequence of length $3^{n+1}$ is constructed in this way. Now, a simple enumeration shows that each different choice of $Z$ and $Y$ yields a distinct SDS. ∎
The same technique and expression presented in Lemma 3 for $\mathbb{Z}_{3}$, can be applied for any alphabet $\mathbb{Z}_{m}$, where $m$ is a prime. For example, if $m=5$, then we start with a SDS Let $S=[X,X+{\bf 1},X+{\bf 2},X+{\bf 3},X+{\bf 4}]$ over $\mathbb{Z}_{5}$, where $X$ is of length $5^{n-1}$. Let $Y_{1}$, $Y_{2}$, and $Y_{3}$ be words of length $5^{n-1}$ and let $Z$ be a word of length $5^{n-1}$ that starts with a zero. From $S$, $Y_{1}$, $Y_{2}$. $Y_{3}$, and $Z$, we form the SDS $V,V+{\bf 1},V+{\bf 2},V+{\bf 3},V+{\bf 4}]$, where
@@ V Non-Binary Single-Track Gray Codes
Let $S_{0},S_{1},\ldots,S_{r-1}$ be $r$ binary self-dual pairwise inequivalent full-order SDSs of length $mn$ over $\mathbb{Z}_{m}$. If for each $i$, $0\leq i<r-1$, $S_{i}$ and $S_{i+1}$ differ in exactly two coordinates, and there also exists an integer $\ell$, where $\text{g.c.d.}(\ell,mn)=1$, such that $S_{r-1}$ and ${\bf E}^{\ell}S_{0}$ differ in exactly $m$ coordinates, then the following words form a length $n$, period mnrmnr STGC.
@@
For $m=3$ and $n=3$, there are three SDSs of length 9 which contain each word of length three over $\mathbb{Z}_{3}$ exactly once. These three sequences are the first three columns of the following array. The other 27 columns of the first three rows form an $3\times 27$ array, which is a length 3 period 27 STGC over $\mathbb{Z}_{3}$. Each column of the $9\times 27$ array after the first three columns is the associated sequence in the required shift. Note that also each of the other rows of this $9\times 27$ array is an SDS.
@@
$\blacksquare$
We consider now another construction of the $m^{m-2}$ SDSs of length $m^{2}$ over $\mathbb{Z}_{m}$. This construction will present us an easier way to form an STGC of length $m$ and period $m^{m}$. Such an STGC will be used as a seed for the recursive construction for an STGC of length $p^{t}$ with period $p^{p^{t}}$, where $p$ is an odd prime. We can order the SDSs of length $m^{2}$ lexicographically such that the first $m$ digits will start with a zero and the sum of these digits will be 0 modulo $m$. This simple ordering contains $m^{m-2}$ sequences of length $m^{2}$. Each word of length $m$ over $\mathbb{Z}_{m}$ is a subsequence of exactly one of these sequences. Now, we use Theorem 8 to form an STGC.
@@
We order this set of differences lexicographically. In this order, there are $m^{m-1}$ sequences of differences of length $m$, where the sum of the differences in each sequence is 1 modulo $m$. Note that two sequences of differences are associated with the same SDS if and only if one is a cyclic shift of the other. From all possible $m$ cyclic shifts, only one sequence of differences is taken. For each sequence of differences, there are $m$ distinct cyclic shifts since the sum 1 modulo $m$ of the differences implies that there is no periodicity in a sequence of differences.
Therefore, there are $m^{m-2}$ cyclic sequences of differences in the final set of differences. This set of sequences containing the differences is ordered in a way that every two consecutive sequences (including the last and the first) have an identical subsequence of length m2m-2. These subsequences imply that in the corresponding SDSs of length $m^{2}$, the $m^{2}$ subsequences of length $m$ can be paired in a way that each pair differs in exactly one coordinate.
@@
For $n=4$ and $m=4$, the 16 sequences of differences are ordered in the following $4\times 16$ array such that every two consecutive columns have an identical subsequence of length 2.
@@
For example we consider $m=3$, and construct the SDS $V,V+{\bf 1},V+{\bf 2}]$ of period $3^{n+1}$, from the SDS $X,X+{\bf 1},X+{\bf 2}]$, where $V=(Z,Z+Y,Z+2Y+X)$, as proved in Lemma 3. The words $Z$ and $Y$ are of length $3^{n-1}$, where $Z$ is any word over $Z_{3}$ that starts with a zero and $Y$ is any word over $\mathbb{Z}_{3}$.
When $Z$ and $Y$ are fixed, then for $X$ we are using the component $X$ from $X,X+{\bf 1},X+{\bf 2}]$ of all the sequences of the STGC of length $3^{n}$ and period $3^{3^{n}}$. The outcome is a component which will form an STGC of length $3^{n+1}$ and period $3^{3^{n}}$.
The next step is more complicated, and in this step, we combine all the components into one maximum period STGC of length $3^{n}$ and period $3^{3^{n}}$. For two components which share the same $Z$ and differ in position $i$ modulo $n$ of $Y$, the $X$’s used will also differ in the $i$-th position in a way that the only part of $V$ in which the two SDSs differ is in one coordinate of $Z+Y$.
@@
The last step is to guarantee in this recursive construction that for each $i$, $1\leq i\leq n$, there are two consecutive words in the list that differ in position $i$. The details for this and the other steps of the long proof of correctness will be presented in the full version of this paper.
The same procedure will work for any given prime $p$, but the merging of the components will become more evolved. For example, when $p=5$ we will have to consider $Z$, $Y_{1}$, $Y_{2}$, and $Y_{3}$ (see the discussion at the end of Section IV). Having the right seeds for $p=3$ and $p=5$, an STGC of length $p^{t}$ and period $p^{p^{t}}$ was constructed for each $t\geq 1$.

### v2

# Tuvi Etzion

@@ Abstract
Binary self-dual sequences have been considered and analyzed throughout the years, and they have been used for various applications. Motivated by a construction for single-track Gray codes, we examine the structure and recursive constructions for binary and non-binary self-dual sequences. The feedback shift registers that generate such sequences are discussed. The connections between these sequences and maximum period single-track codes are discussed. Maximum period non-binary single-track Gray codes of length $p^{t}$ and period $p^{p^{t}}$ are constructed. These are the first infinite families of maximum period codes presented in the literature.
@@ I Introduction
Motivated by these applications and mainly by single-track Gray codes, we are trying to find properties of these sequences, to construct them recursively and efficiently, and to order them for the construction of single-track Gray codes. The Gray code, which are considered, will be only for maximum period codes. This will restrict some of the parameters of the self-dual sequences. We consider the generalization of such sequences over a non-binary alphabet, where a non-binary self-dual sequence is a sequence that is invariant under the addition of a nonzero constant to all the elements of the sequence.
@@ II Preliminaries
The consecutive values obtained from $x_{n+1}$ (until the initial state is repeated), starting with any initial state is the sequence generated by the register. The CCRn generates all the SDSs whose period is a divisor of $2n$ and no other sequences, where the period of a sequence $S=[s_{1}s_{2}~\cdots~s_{k}]$ is the length of the shortest word $X$ for which $S=[X,X,\ldots,X]$. If the FSRn is linear, then it can also be represented by a polynomial of degree $n$ over the binary field.
The following enumeration was presented in [11].
@@
Let SD(n)(n) denote the number of SDSs of period $n$. A connection between the number of sequences generated by the CCRn and the number of SDSs was proved in [6].
@@
where $Z$ can be any sequence of length $2^{n}$ (note that plus and minus are the same in the binary field). There are $2^{2^{n}-1}$ distinct cyclic sequences generated in this way (if $Z$ is replaced by $Z+X+Y$ the same sequence is obtained). If $Y=\bar{X}$ then
@@
It was proved in [17] that there is no maximum period STGC of length $2^{t}$ and period $2^{2^{t}}$, $t>1$. An STGC of length $2^{t}$ and period $2^{2^{t}}-2^{t+1}$ was constructed in [9]. Non-binary STGCs of length $p^{m}$ and period $p^{p^{m}}$ are discussed in Section V. These are the only possible maximum period STGCs that contain all the words of the given length. For each odd prime $p$, we might have a maximum period STGC of length $p$ and period $2^{p}-2$. Such codes are known to exist for all primes up to $p=19$ [4], and they are constructed using the construction of Theorem 3, but they can also be obtained by the construction of Theorem 4. Maximum period STGC using these theorems can also be obtained for $n=9$ and $n=25$. For other values of $n$, the construction of Theorem 3 cannot yield a maximum period STGCs. The situation for the construction of Theorem 4 is quite different. If $n=2p$, then there is exactly one SDS of period 4 and $\frac{2^{2p}-4}{4p}$ SDSs of period $4p$ generated by the CCR2p. It might be possible, in this case, to use the construction of Theorem 4 to generate a maximum period STGC. If $n=4p$, then there are exactly two SDSs of period 8 and $\frac{2^{4p}-16}{8p}$ SDSs of period $8p$ generated by the CCR4p. It might be possible, in this case, to use the construction of Theorem 4 to generate a maximum period STGC. Such analysis can be further given. If $n=8p$, then there are exactly sixteen SDSs of period 16 and $\frac{2^{8p}-256}{16p}$ SDSs of period $16p$ generated by the CCR8p and it might be possible to use the construction of Theorem 4 to generate a maximum period STGC when $p>16$. The same process can be analyzed for $n=2^{i}p$, $i\geq 1$.
@@ III Binary Self-Dual Sequences
The operation ${\bf D}^{2^{n}}=({\bf E}+1)^{2^{n}}$ and its inverse are associated with the sequences generated by the polynomial $(x+1)^{2^{n}}$. They are generalized for binary sequences whose length is not a power of 2 in the following way. Let ${\bf\Delta}_{m}$ be the following operator defined on binary sequences whose length is $rm$. Let $S=[X_{1},X_{2},\ldots,X_{r}]$ be a binary sequence, where $X_{i}$, $1\leq i\leq r$, is a word of length of length $m$. The operator ${\bf\Delta}_{m}$ is defined on $S$ as
@@
If $\cal S$} is a set of sequences, then we define ${\bf\Delta}_{m}{\cal S}$ to be the set $\{{\bf\Delta}_{m}T~:~T\in{\cal S}\}$. and ${\bf\Delta}_{m}^{-1}{\cal S}$ is the set $\{{\bf\Delta}_{m}^{-1}T~:~T\in{\cal S}\}$. It is readily verified that for any sequence $S$ of length $rm$ we have that ${\bf\Delta}_{m}({\bf\Delta}_{m}^{-1}S)={\bf\Delta}^{-1}_{m}({\bf\Delta}_{m}S)=S$.
@@
An SDS of length $2n$ generated by the CCRn can be written in the form $X,\bar{X}]$, where the length of $X$ is $n$. Hence, ${\bf\Delta}_{n}^{-1}[X,\bar{X}]=[Y,X+Y,\bar{Y},X+\bar{Y}]$, where $Y$ is any word of length $n$, has length $4n$ and it is an SDS generated by the CCR2n. ∎
@@
There are two SDSs generated by the CCR3, $[000111]$ and $[01]$. By applying ${\bf\Delta}_{3}^{-1}$ on these two sequences, the six sequences of the CCR6 are obtained. Four sequences
@@
of period 12 are obtained by applying ${\bf\Delta}_{3}^{-1}$ on $[000111]$. One sequence of period 12, [000010111101][000~~010~~111~~101], and one sequence of period 4, $[0011]$, are obtained by applying ${\bf\Delta}_{3}^{-1}$ on $[01]$.
We can continue and apply ${\bf\Delta}_{6}^{-1}$ on these 6 SDSs of the CCR6. By applying ${\bf\Delta}_{6}^{-1}$ on the five SDSs of period 12, we obtain 160 SDSs of period 24, which are generated by the CCR12. By applying ${\bf\Delta}_{6}^{-1}$ on the sequence $[0011]$ we obtain the two SDSs $[00001111]$ and $[00101101]$ of period 8 and ten SDSs of period 24, all generated by the CCR12.  $\blacksquare$
@@
The recurrence of Theorem 5 and the enumeration of Theorem 6 and Example 1 can be further generalized and analyzed. From the $\frac{2^{2^{i+1}p}-2^{2^{i+1}}}{2^{i+2}p}$ SDSs of period $2^{i+2}p$ only $\frac{2^{2^{i}(p+1)}-2^{2^{i+1}}}{2^{i+2}p}$ are generated from SDSs of period $2^{i+1}$, where each such SDS of period $2^{i+2}p$ can be generated using any one of $p$ different words of length $2^{i}p-1$. It can be analyzed which words of length $2^{i}p-1$ yield SDSs of period $2^{i+2}$ and which yield SDSs of period $2^{i+2}p$.
@@ IV Non-binary Self-Dual Sequences
Linear recurring sequences over rings were extensively studied, but the associated shift registers were hardly investigated (see [13, 18] and related references) and the same is for non-binary self-dual sequences. The definition of the CCRn is genralized for an FSRn over $\mathbb{Z}_{m}$. The $m$-CCRn is an FSRn over $\mathbb{Z}_{m}$ with feedback function
@@
where $x_{i}\in\mathbb{Z}_{m}$, $1\leq i\leq n$, and the computation is performed modulo $m$. The function can also be taken as $f(x_{1},x_{2},\ldots,x_{n})=x_{1}+r$, where $r$ is relatively prime to $m$. The enumerations of the number of sequences of the CCRn (see Theorem 1 and Lemma 1) can be generalized.
@@
$\sum_{\begin{subarray}{c}d|mn,~d\nmid n\\ d|n-\frac{d}{m}\end{subarray}}\textup{SD}_{m}(d).$
@@
The sequence $V,V+{\bf 1},V+{\bf 2}]$ (a bold integer is a sequence with only this integer), where $V=(Z,Z+Y,Z+2Y+X)$, is an SDS of period $3^{n+1}$ over $\mathbb{Z}_{3}$. Each different choice of $Z$ and $Y$ yields a distinct SDS. Each SDS of period $3^{n+1}$ over $\mathbb{Z}_{3}$ is constructed in this way.
@@
The proof has similarity to that of Theorem 5 with one significant exception. By the definition of ${\bf D}^{-1}$ we have D^-3^n-1 [X , X+1, X+2] = [Y,Y+X, Y+2X+ 1], D^-3^n-1 [Y,Y+X, Y+2X+ 1] = [V , V+1, V+2], where $V=(Z,Z+Y,Z+2Y+X)$. Note, that since either $V$ or $V+{\bf 1}$ or $V+{\bf 2}$ starts with a zero, it was sufficient to consider for $Z$ only those words of length $3^{n-1}$ that start with a zero.
@@
This implies that each SDS of length $3^{n+1}$ is constructed in this way. Now, a simple enumeration shows that each different choice of $Z$ and $Y$ yields a distinct SDS. ∎
The same technique and expression presented in Lemma 3 for $\mathbb{Z}_{3}$, can be applied for any alphabet $\mathbb{Z}_{m}$, where $m$ is a prime. For example, if $m=5$, then we start with a SDS Let $S=[X,X+{\bf 1},X+{\bf 2},X+{\bf 3},X+{\bf 4}]$ over $\mathbb{Z}_{5}$, where $X$ is a word of length $5^{n-1}$. Let $Y_{1}$, $Y_{2}$, and $Y_{3}$ be words of length $5^{n-1}$ and let $Z$ be a word of length $5^{n-1}$ that starts with a zero. From $S$, $Y_{1}$, $Y_{2}$. $Y_{3}$, and $Z$, we form the SDS $V,V+{\bf 1},V+{\bf 2},V+{\bf 3},V+{\bf 4}]$, where
@@ V Non-Binary Single-Track Gray Codes
Let $S_{0},S_{1},\ldots,S_{r-1}$ be $r$ binary self-dual pairwise inequivalent full-order SDSs of length $mn$ over $\mathbb{Z}_{m}$. If for each $i$, $0\leq i<r-1$, $S_{i}$ and $S_{i+1}$ differ in exactly $m$ coordinates, and there also exists an integer $\ell$, where $\text{g.c.d.}(\ell,mn)=1$, such that $S_{r-1}$ and ${\bf E}^{\ell}S_{0}$ differ in exactly $m$ coordinates, then the following words form a length $n$, period mnrmnr STGC.
@@
For $m=3$ and $n=3$, there are three SDSs of length 9 which contain each word of length three over $\mathbb{Z}_{3}$ exactly once. These three sequences are the first three columns of the following array. The other 27 columns of the first three rows form an $3\times 27$ array, which is a length 3 period 27 STGC over $\mathbb{Z}_{3}$. Each column of the $9\times 27$ array after the first three columns is the associated sequence in the required shift. Note that also each of the rows of this $9\times 27$ array is an SDS.
@@
We consider now a construction of the $m^{m-2}$ SDSs of length $m^{2}$ over $\mathbb{Z}_{m}$, which is an easier way to form an STGC of length $m$ and period $m^{m}$. Such an STGC will be used as a seed for the recursive construction for an STGC of length $p^{t}$ with period $p^{p^{t}}$, where $p$ is an odd prime. We can order the SDSs of length $m^{2}$ lexicographically such that the first $m$ digits will start with a zero and the sum of these digits will be 0 modulo $m$. This simple ordering contains $m^{m-2}$ sequences of length $m^{2}$. Each word of length $m$ over $\mathbb{Z}_{m}$ is a subsequence of exactly one of these sequences. Now, we use Theorem 8 to form an STGC.
@@
We order this set of differences lexicographically. In this order, there are $m^{m-1}$ sequences of differences of length $m$, where the sum of the differences in each sequence is 1 modulo $m$. Two sequences of differences are associated with the same SDS if and only if one is a cyclic shift of the other. From all possible $m$ cyclic shifts, only one sequence of differences is taken. For each sequence of differences, there are $m$ distinct cyclic shifts since the sum 1 modulo $m$ of the differences implies that there is no periodicity in a sequence of differences.
Therefore, there are $m^{m-2}$ cyclic sequences of differences in the final set of differences. This set of sequences containing the differences is ordered in a way that every two consecutive sequences (including the last and the first) have an identical subsequence of length m2m-2. Therefore, in the corresponding SDSs of length $m^{2}$, the $m^{2}$ subsequences of length $m$ can be paired such that each pair differs in exactly one coordinate.
@@
For $n=m=4$, the 16 sequences of differences are ordered in a $4\times 16$ array such that every two consecutive columns have an identical subsequence of length 2.
@@
For example we consider $m=3$, and construct the SDS $V,V+{\bf 1},V+{\bf 2}]$ of period $3^{n+1}$, from the SDS $X,X+{\bf 1},X+{\bf 2}]$, where $V=(Z,Z+Y,Z+2Y+X)$, as proved in Lemma 3. The words $Z$ and $Y$ are of length $3^{n-1}$, where $Z$ is any word over $Z_{3}$ that starts with a zero and $Y$ is any word of length $3^{n-1}$ over $\mathbb{Z}_{3}$.
When $Z$ and $Y$ are fixed, then for $X$ we are using the component $X$ from $X,X+{\bf 1},X+{\bf 2}]$ of all the sequences of the STGC of length $3^{n}$ and period $3^{3^{n}}$. The outcome is a component yielding an STGC of length $3^{n+1}$ and period $3^{3^{n}}$.
The next step is more complicated, and in this step, we merge all the components into one maximum period STGC of length $3^{n+1}$ and period $3^{3^{n+1}}$. For two components which share the same $Z$ and differ in position $i$ modulo $n$ of $Y$, the $X$’s used will also differ in the $i$-th position in a way that the only part of $V$ in which the two SDSs differ is in one coordinate (the $i$-th coordinate) of $Z+Y$.
@@
The last step is to guarantee in this recursive construction that for each $i$, $1\leq i\leq n$, there are two consecutive words in the list that differ in position $i$. Details will be presented in the full version of this paper.
The same procedure will work for any given prime $p$, but merging the components will become more evolved. For example, when $p=5$ we will have to consider $Z$, $Y_{1}$, $Y_{2}$, and $Y_{3}$ (see the discussion at Section IV). Having the right seeds for $p=3$ and $p=5$ (which were easily obtained), an STGC of length $p^{t}$ and period $p^{p^{t}}$ was constructed for each $t\geq 1$.